f(x + h) – f(x) Simplify the difference quotient for the given function h f(x) = 6x2 - 5x + 5 f(x+h)-0 f(x+h)-f(x) = h (SimplPlease show steps, I'm rather lost.
Answer #1

f(x) = 6x2 sx +s, f/2th) = 6/2+b]²5/4th) +5 fluth)-f(x) = 6/9th) 2_s/ath)+5 - 6x² +52-5 6/2²+ 2ah the) - 5h-6a2 h (12x+6h-5)